# Robust Clustering on High-Dimensional Data with Stochastic Quantization

### v1

@@ 4 Data Encoding with Triplet Network
Recent research by [35] and [54] has employed a Triplet Network architecture to learn features from high-dimensional discrete image data and encode them into low-dimensional representations in the latent space. These authors proposed a semi-supervised learning approach where the Triplet Network is trained on a labeled subset of data to encode them into latent representations in $\mathbb{R}^{n}$, and subsequently used to project the remaining unlabeled fraction onto the same latent space. This approach significantly reduces the time and labor required for data annotation without compromising accuracy.
@@
$\mathcal{L}_{triplet}=\max(0,d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{+}_{i}))-d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{-}_{i}))+\alpha)$ (34)
@@
In our research, we utilized a Convolutional Network architecture as $f_{\theta}$, as proposed by [31]. The detailed overview of the architecture, its training using the Backpropagation algorithm, and accuracy evaluation are beyond the scope of this paper; however, [5], [28], and [31] provide extensive coverage of these topics.
Regarding triplet mining strategies for $(\xi_{i},\xi^{+}_{i},\xi^{-}_{i})$, it is crucial to select an approach that produces the most informative gradients for the objective function (34). [59] discusses various online triplet mining strategies, which select triplets within each batch of a training set on each iteration. We employed the semi-hard triplet mining strategy, which chooses an anchor-negative pair that is farther than the anchor-positive pair but within the margin $\alpha$:
$\xi^{-}_{i}=\operatorname*{arg\,min}_{\begin{subarray}{c}\xi:C(\xi_{i})\neq C(\xi)\\ d(f_{\theta}(\xi_{i}),f_{\theta}(\xi))>d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{+}_{i}))\end{subarray}}d(f_{\theta}(\xi_{i}),f_{\theta}(\xi))$ (35)
@@ 5 Numerical Experiments
We approached the image classification task as a semi-supervised learning problem, training models on varying fractions of labeled training data (10%, 30%, 60%, 80%, and 100%). The training dataset was split using uniform sampling into labeled and unlabeled portions according to the specified percentages. For the Triplet Network, we employed a Convolutional Neural Network architecture consisting of two Convolutional Layers with 3x3 filters (feature map dimensions of 32 and 64, respectively, with $\text{stride}=1$ and $\text{padding}=1$), followed by 2x2 Max-Pooling layers, and two Dense Layers. ReLU activation functions were used throughout the network. We trained separate Triplet Network models for each labeled data fraction with the following hyperparameters: 50 epochs, batch size of $1000\times\text{fraction}$, learning rate $\rho=10^{-3}$, and $l_{2}$ regularization rate $\lambda=10^{-5}$. For the triplet loss (34) and triplet mining (35), we set the margin hyperparameter $\alpha=1.0$. To facilitate meaningful feature capture while enabling visualization, we chose a latent space dimensionality of $\mathbb{R}^{3}$ (see Fig. 3).
@@
The Triplet Network was used to project latent representations onto $\mathbb{R}^{3}$ space from both labeled and unlabeled training data. These representations were then used to train the Stochastic Quantization algorithm as an unsupervised learning model. For each set of latent representations corresponding to different labeled data fractions, we trained the Stochastic Quantization algorithm and its adaptive variants (16-2.1). We employed the KMeans++ initialization strategy (25) for all variants, with a rank hyperparameter $r=3$. To ensure convergence, we used different learning rates for each variant: $\rho=0.001$ for SGD, Momentum, and NAG; $\rho=0.9$ for AdaGrad; and $\rho=0.01$ for RMSProp and ADAM. With these hyperparameters, all Stochastic Quantization variants converged to the global optima (see Fig. 4), with most converging on the first iteration (see Fig. 5).

### v2

@@ 4 Data Encoding with Triplet Network
Recent research [34, 54] has employed a Triplet Network architecture to learn features from high-dimensional discrete image data and encode them into low-dimensional representations in the latent space. The authors proposed a semi-supervised learning approach where the Triplet Network is trained on a labeled subset of data to encode them into latent representations in $\mathbb{R}^{n}$, and subsequently used to project the remaining unlabeled fraction onto the same latent space. This approach significantly reduces the time and labor required for data annotation without compromising accuracy.
@@
$\mathcal{L}_{triplet}(\theta)=\max(0,d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{+}_{i}))-d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{-}_{i}))+\alpha)$ (34)
@@
In our research, we utilized a Convolutional Network architecture as $f_{\theta}$, as proposed by [30]. The detailed overview of the architecture, its training using the Backpropagation algorithm, and accuracy evaluation are beyond the scope of this paper; [5], [27], and [30] provide extensive coverage of these topics. However, notice that function (34) is non-smooth and non-convex and the standard Backpropagation technique is not validated for such case. The extension of this technique to the non-smooth non-convex case was made in [39].
Regarding triplet mining strategies for $(\xi_{i},\xi^{+}_{i},\xi^{-}_{i})$, it is crucial to select an approach that produces the most informative gradients for the objective function (34). Paper [59] discusses various online triplet mining strategies, which select triplets within each batch of a training set on each iteration. We employed the semi-hard triplet mining strategy, which chooses an anchor-negative pair that is farther than the anchor-positive pair but within the margin $\alpha$:
$\xi^{-}_{i}=\operatorname*{arg\,min}_{\begin{subarray}{c}\xi:C(\xi)\neq C(\xi_{i})\\ d(f_{\theta}(\xi_{i}),f_{\theta}(\xi))>d(f_{\theta}(\xi_{i}),f_{\theta}(\xi^{+}_{i}))\end{subarray}}d(f_{\theta}(\xi_{i}),f_{\theta}(\xi))$ (35)
@@ 5 Numerical Experiments
We approached the image classification task as a semi-supervised learning problem, training models on varying fractions of labeled training data (10%, 30%, 60%, 80%, and 100%). The training dataset was split using uniform sampling into labeled and unlabeled portions according to the specified percentages. For the Triplet Network, we employed a Convolutional Neural Network architecture consisting of two Convolutional Layers with 3x3 filters (feature map dimensions of 32 and 64, respectively, with $\text{stride}=1$ and $\text{padding}=1$), followed by 2x2 Max-Pooling layers, and two Dense Layers. ReLU (non-differentiable) activation functions were used throughout the network. Together with non-smooth Triplet loss function this makes the learning problem highly non-smooth and non-convex (see discussion of this issue in [39]). We trained separate Triplet Network models for each labeled data fraction with the following hyperparameters: 50 epochs, batch size of $1000\times\text{fraction}$, learning rate $\rho=10^{-3}$, and $l_{2}$ regularization rate $\lambda=10^{-5}$. For the triplet loss (34) and triplet mining (35), we set the margin hyperparameter $\alpha=1.0$. To facilitate meaningful feature capture while enabling visualization, we chose a latent space dimensionality of $\mathbb{R}^{3}$ (see Fig. 3).
@@
The Triplet Network was used to project latent representations onto $\mathbb{R}^{3}$ space from both labeled and unlabeled training data. These representations were then used to train the Stochastic Quantization algorithm as an unsupervised learning model. For each set of latent representations corresponding to different labeled data fractions, we trained the Stochastic Quantization algorithm and its adaptive variants (16)–(2.1) from subsection 2.1. We employed the K-Means++ initialization strategy (25) for all variants, with a rank hyperparameter $r=3$. To ensure convergence, we used different learning rates for each variant: $\rho=0.001$ for SGD, Momentum, and NAG; $\rho=0.9$ for AdaGrad; and $\rho=0.01$ for RMSProp and ADAM. With these hyperparameters, all Stochastic Quantization variants converged to the global optima (see Fig. 4), with most converging on the first iteration (see Fig. 5).